OpenStudy (anonymous):

find the x and y-intercept of each line y=7x+5 please show step by step thank you find the x and y-intercept of each line y=7x+5 please show step by step thank you @Mathematics

jimthompson5910 (jim_thompson5910):

To find the x-intercept, plug in y=0 and solve for x y=7x+5 0=7x+5 -5=7x -5/7 = x x=-5/7 So the x-intercept is (-5/7, 0) ------------------- To find the y-intercept, plug in x=0 and solve for y y=7x+5 y=7(0)+5 y=0+5 y=5 So the y-intercept is (0, 5)

OpenStudy (anonymous):

Thanks!! <3